OpenStudy (anonymous):

Help PLease!!! Joshua Bennett purchased a television for $245.95 with 40.25% down. He will make 6 payments of $25.83 each. What is the total amount paid (will include down payment) and finance charge in dollars? The total paid (including down payment) is $__. 253.97 253.47 252.97 The finance charge in dollars is $___. 7.02 7.52 8.02

OpenStudy (anonymous):

Well to find the down payment you can do proportions \[\frac{ x }{ 245.95 }=\frac{ 40.25 }{ 100 }\]Then turns to\[100x=9899.4875\]Then x equals to\[x=98.99\] so that's Joshua's down payment. Next to find what else he had to pay, it says he paid 25.83 six times. So\[6*25.83=154.98\]Then add 98.99 and 154.98 together to get Joshua's total paid

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Yep hope it helps and Good Luck :)